OpenStudy (anonymous):

is x+1/x^2-1 equivalent to x/x-1

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{x+1}{(x+1)(x-1)}=\frac{1}{x-1}\] so long as x is not -1

OpenStudy (anonymous):

is 2x+4/4 equivalent to 2

OpenStudy (angela210793):

two equations are equivalent when they have the same solution set (if i remember correctly)

OpenStudy (anonymous):

I would have said, thanks, but what ever

OpenStudy (anonymous):

no it isn't

OpenStudy (anonymous):

\[\frac{2x+4}{4}=\frac{2(x+2)}{4}=\frac{x+2}{2}\]

OpenStudy (anonymous):

is x^2y^3z/xy^2 equivanent to xz/y^-1